# Assessing and improving reliability of neighbor embedding methods: a map-continuity perspective

### v1

@@ E.3.2 Hyperbolic Structure in the LOO Loss: Proof of Theorem 1
$-\nabla_{{\bm{y}}}L({\bm{y}};{\bm{x}}^{\varepsilon})=\bm{F}_{a}+\bm{F}_{r},\hskip 18.49988pt\hskip 18.49988pt\text{where}$
$\bm{F}_{a}=4\sum_{i=1}^{n}\frac{v_{i,n+1}}{1+\|{\bm{y}}_{i}-{\bm{y}}\|^{2}}({\bm{y}}_{i}-{\bm{y}}),$
$\bm{F}_{r}=-\frac{4}{Z}\sum_{i=1}^{n}\frac{1}{\big(1+\|{\bm{y}}_{i}-{\bm{y}}\|^{2}\big)^{2}}({\bm{y}}_{i}-{\bm{y}})$
Let us simplify $\bm{F}_{a}$ and $\bm{F}_{r}$ using the asymptotics we assumed in the theorem. First we observe that
$\frac{1}{1+\|\pm{\bm{\theta}}-{\bm{y}}+\bm{\delta}_{i}\|^{2}}$ $=\frac{1}{1+\|{\bm{\theta}}\|^{2}+2\langle\pm{\bm{\theta}},-{\bm{y}}+\bm{\delta}_{i}\rangle+\|-{\bm{y}}+\bm{\delta}_{i}\|^{2}}$
$=\frac{1}{\|{\bm{\theta}}\|^{2}}\left[1-2\Big\langle\pm\frac{{\bm{\theta}}}{\|{\bm{\theta}}\|^{2}},-{\bm{y}}+\bm{\delta}_{i}\Big\rangle+O\Big(\frac{1}{\|{\bm{\theta}}\|^{2}}\Big)\right]$
@@
$\frac{{\bm{y}}_{i}-{\bm{y}}}{1+\|{\bm{\theta}}-{\bm{y}}+\bm{\delta}_{i}\|^{2}}$ $=\frac{{\bm{\theta}}+\bm{\delta}_{i}-{\bm{y}}}{\|{\bm{\theta}}\|^{2}}\left[1-2\Big\langle\frac{{\bm{\theta}}}{\|{\bm{\theta}}\|^{2}},-{\bm{y}}+\bm{\delta}_{i}\Big\rangle+O\Big(\frac{1}{\|{\bm{\theta}}\|^{2}}\Big)\right]$
$=\frac{{\bm{\theta}}}{\|{\bm{\theta}}\|^{2}}+\frac{\bm{\delta}_{i}-{\bm{y}}}{\|{\bm{\theta}}\|^{2}}-\frac{2{\bm{\theta}}}{\|{\bm{\theta}}\|^{2}}\Big\langle\frac{{\bm{\theta}}}{\|{\bm{\theta}}\|^{2}},-{\bm{y}}+\bm{\delta}_{i}\Big\rangle+O\Big(\frac{1}{\|{\bm{\theta}}\|^{3}}\Big);$
@@
$\frac{{\bm{y}}_{i}-{\bm{y}}}{1+\|-{\bm{\theta}}-{\bm{y}}+\bm{\delta}_{i}\|^{2}}$ $=\frac{-{\bm{\theta}}+\bm{\delta}_{i}-{\bm{y}}}{\|{\bm{\theta}}\|^{2}}\left[1-2\Big\langle-\frac{{\bm{\theta}}}{\|{\bm{\theta}}\|^{2}},-{\bm{y}}+\bm{\delta}_{i}\Big\rangle+O\Big(\frac{1}{\|{\bm{\theta}}\|^{2}}\Big)\right]$
$=-\frac{{\bm{\theta}}}{\|{\bm{\theta}}\|^{2}}+\frac{\bm{\delta}_{i}-{\bm{y}}}{\|{\bm{\theta}}\|^{2}}-\frac{2{\bm{\theta}}}{\|{\bm{\theta}}\|^{2}}\Big\langle\frac{{\bm{\theta}}}{\|{\bm{\theta}}\|^{2}},-{\bm{y}}+\bm{\delta}_{i}\Big\rangle+O\Big(\frac{1}{\|{\bm{\theta}}\|^{3}}\Big)\,.$
Since $\sum_{i}\bm{\delta}_{i}=\mathbf{0}$, we have
$\frac{1}{n}\sum_{i=1}^{n}\frac{{\bm{y}}_{i}-{\bm{y}}}{1+\|{\bm{y}}_{i}-{\bm{y}}\|^{2}}=-\frac{{\bm{y}}}{\|{\bm{\theta}}\|^{2}}+\frac{2}{\|{\bm{\theta}}\|^{2}}{\bm{y}}_{\mathbin{\!/\mkern-5.0mu/\!}}+O\Big(\frac{1}{\|{\bm{\theta}}\|^{3}}\Big)\,.$
@@
$\frac{1}{n}\sum_{i\in{\mathcal{I}}_{+}}\frac{{\bm{y}}_{i}-{\bm{y}}}{1+\|{\bm{y}}_{i}-{\bm{y}}\|^{2}}-\frac{1}{n}\sum_{i\in{\mathcal{I}}_{-}}\frac{{\bm{y}}_{i}-{\bm{y}}}{1+\|{\bm{y}}_{i}-{\bm{y}}\|^{2}}=\frac{2{\bm{\theta}}}{\|{\bm{\theta}}\|^{2}}+O\Big(\frac{1}{\|{\bm{\theta}}\|^{3}}\Big)\,.$
@@
$\bm{F}_{a}=\frac{4np_{0}(2{\bm{y}}_{\mathbin{\!/\mkern-5.0mu/\!}}-{\bm{y}})}{\|{\bm{\theta}}\|^{2}}+\frac{8n\epsilon{\bm{\theta}}}{\|{\bm{\theta}}\|^{2}}+O\Big(\frac{1}{\|{\bm{\theta}}\|^{3}}\Big).$ (S5)
To handle $\bm{F}_{r}$, we notice that
$\frac{{\bm{y}}_{i}-{\bm{y}}}{\big(1+\|{\bm{y}}_{i}-{\bm{y}}\|^{2}\big)^{2}}=O\Big(\frac{1}{\|{\bm{\theta}}\|^{3}}\Big)\,.$
@@
$\frac{1}{1+\|{\bm{y}}_{i}-{\bm{y}}_{j}\|^{2}}\leq 1,\qquad\Longrightarrow\qquad Z\leq n(n+1)\,.$
@@
$Z\geq\sum_{i,j\in{\mathcal{I}}_{+}}\frac{1}{1+\|{\bm{y}}_{i}-{\bm{y}}_{j}\|^{2}}=\sum_{i,j\in{\mathcal{I}}_{+}}\frac{1}{1+\|\bm{\delta}_{i}-\bm{\delta}_{j}\|^{2}},.$
@@
$\bm{F}_{r}=O\Big(\frac{1}{\|{\bm{\theta}}\|^{3}}\Big)\,.$ (S6)
Combining Eqn. S5 and S6, we reach our conclusion.
E.4 Supplementary Tables and Figures
@@
Table S1: Datasets analyzed in this work. This paper analyzes six real-world datasets: CIFAR-10, Describable Textures Dataset (DTD), the mouse embryonic stem cell differentiation data (Embryo), the single-cell RNA-seq dataset generated from PBMCs treated with interferon-$\beta$ (IFNB), the mouse brain single-cell ATAC-seq data (Brain), the single-cell RNA-seq dataset generated from human pancreatic tissues (Panc8). The table lists the number of points, dimensions, domains, and usage of the datasets.
@@
(a) Tests of singularity score pattern in single-cell RNA-seq data of mouse embryonic stem cells (mESCs) differentiation.
@@
(b) Tests of singularity score pattern in mouse brain single-cell ATAC-seq data.
Table S4: Tests of the distribution difference of singularity scores in (a) mouse embryonic stem cell differentiation data and (b) mouse brain chromatin accessibility data. We verifiy the distribution difference using Spearman’s rank correlation tests between singularity scores and distances to cluster centers, F-tests and per- mutation tests for a local regression model (singularity scores regressed against locations). The results of all three tests confirm the distribution difference of singularity scores between small and large perplexities.
@@
Figure S2: Large perplexity lessens FI discontinuity. We use the same mixture data as in Figure 3 and run the standard t-SNE algorithm 6 times at perplexity 5, 20, 40, 60, 80, 100. a Contour plot shows the landscape of LOO loss $L({\bm{y}};{\bm{x}})$ for ${\bm{x}}=0.5({\bm{c}}_{1}+{\bm{c}}_{2})$ under different perplexities. The number of local minima in the loss decreases with higher perplexity, indicating that the FI discontinuity lessens under higher perplexity. b Trajectories of the embedding point of $\bm{x}$} are shown with different perplexities. Under small perplexities, numerous local minima cause an uneven trajectory of embedding points when we add $\bm{x}$} at evenly interpolated locations; while under larger perplexities, the trajectory is more smooth. This further suggests a reduction of FI discontinuity when the perplexity increases. c Embeddings with dichotomized singularity scores. Embedding points with high singularity scores decrease in number when increasing the perplexity.
@@
Figure S3: Comparing loss landscapes for single-cell data with different perplexities. We present LOO loss landscapes under small and large perplexities for mouse embryonic stem cells differentiation data (a-b) and mouse brain single-cell ATAC-seq data (c-d). We randomly choose an input ${\bm{x}}_{i}$ from the dataset and plot the landscape of the partial LOO loss $L_{i}({\bm{y}};{\bm{x}}_{i})$. We observe that more local minima emerge at random locations in the partial LOO loss landscape under a small perplexity than under a large perplexity, which supports our claim about the reduction of local minima and FI discontinuity under a large perplexity.
@@
Figure S4: Perturbation scores of CIFAR-10 image features with different choices of perturbation length. We vary the choice of the perturbation length $\lambda\in\{1,2,3\}$ in the calculation of perturbation scores; see Eqn. 6. We find that embedding points receiving high scores are consistent across different choices of $\lambda$. This suggests that perturbation scores are not sensitive to the choice of the perturbation length.
@@
Figure S5: Pre-screening points for calculating perturbations scores provides comparable results. We examine the validity of the pre-screening step by comparing the perturbation scores with the pre-screening step (visualized in left panels) and those without (visualized in right panels). We present the results of perturbation scores on three datasets—5-component Gaussian mixture data (a), Swiss roll data (b), and deep learning feature data of CIFAR-10 (c). We find that the calculating perturbation scores with the pre-screening step still identifies most of the OI discontinuity locations, so pre-screening provides a faster and comparably reliable assessment of OI discontinuity.
@@
Figure S6: The approximation method 1&2 are accurate for calculating perturbation scores. We test the approximation quality on a dataset consisting of 5000 deep learning feature vectors obtained from CIFAR-10 images. a Scatter plot of exact perturbation scores v.s. perturbation scores by the approximation method 1. b Scatter plot of exact perturbation scores v.s. perturbation scores by approximation method 2. We find that perturbation scores using both approximation methods are approximately equal to the exact perturbation scores for almost all the points.

### v2

@@ A.3.2 Hyperbolic Structure in the LOO Loss: Proof of Theorem 1
$-\nabla_{{\mathbf{y}}}L({\mathbf{y}};{\mathbf{x}}^{\varepsilon})=\mathbf{F}_{a}+\mathbf{F}_{r},\qquad\qquad\text{where}$
$\mathbf{F}_{a}=4\sum_{i=1}^{n}\frac{v_{i,n+1}}{1+\|{\mathbf{y}}_{i}-{\mathbf{y}}\|^{2}}({\mathbf{y}}_{i}-{\mathbf{y}}),$
$\mathbf{F}_{r}=-\frac{4}{Z}\sum_{i=1}^{n}\frac{1}{\big(1+\|{\mathbf{y}}_{i}-{\mathbf{y}}\|^{2}\big)^{2}}({\mathbf{y}}_{i}-{\mathbf{y}})$
Let us simplify $\mathbf{F}_{a}$ and $\mathbf{F}_{r}$ using the asymptotics we assumed in the theorem. First we observe that
[…]
@@
[…]
@@
[…]
Since $\sum_{i}{\mathbf{\updelta}}_{i}=\mathbf{0}$, we have
[…]
@@
[…]
@@
[…]
To handle $\mathbf{F}_{r}$, we notice that
[…]
@@
$\frac{1}{1+\|{\mathbf{y}}_{i}-{\mathbf{y}}_{j}\|^{2}}\leq 1,\qquad\Longrightarrow\qquad Z\leq n(n+1)\,.$
@@
$Z\geq\sum_{i,j\in{\mathcal{I}}_{+}}\frac{1}{1+\|{\mathbf{y}}_{i}-{\mathbf{y}}_{j}\|^{2}}=\sum_{i,j\in{\mathcal{I}}_{+}}\frac{1}{1+\|{\mathbf{\updelta}}_{i}-{\mathbf{\updelta}}_{j}\|^{2}},.$
@@
[…]
Combining Eqn. S6 and S7, we reach our conclusion.
A.4 Supplementary Tables and Figures
@@
Mammary [55] 25,806 2,000 Single-cell Singularity score
Table S1: Datasets analyzed in this work. This paper analyzes seven real-world datasets: CIFAR-10, Describable Textures Dataset (DTD), the mouse embryonic stem cell differentiation data (Embryo), the single-cell RNA-seq dataset generated from PBMCs treated with interferon-$\beta$ (IFNB), the mouse brain single-cell ATAC-seq data (Brain), the single-cell RNA-seq dataset generated from human pancreatic tissues (Panc8), the single-cell dataset of mammary epithelial cells (Mammary). The table lists the number of points, dimensions, domains, and usage of the datasets.
@@
Table S4: Tests of the distribution difference of singularity scores in the single-cell RNA-seq data of mouse embryonic stem cells differentiation. We verifiy the distribution difference using Spearman’s rank correlation tests between singularity scores and distances to cluster centers, F-tests and per- mutation tests for a local regression model (singularity scores regressed against locations). The results of all three tests confirm the distribution difference of singularity scores between small and large perplexities.
@@
Table S5: Tests of the distribution difference of singularity scores in the mouse brain single-cell ATAC-seq data. We verifiy the distribution difference using Spearman’s rank correlation tests between singularity scores and distances to cluster centers, F-tests and per- mutation tests for a local regression model (singularity scores regressed against locations). The results of all three tests confirm the distribution difference of singularity scores between small and large perplexities.
Perplexity 30 Perplexity 175
NP1 $0.9962$ $1.0000$ $0.3623$ $1.0000$
NP2 $0.0557$ $0.2223$ $0.4702$ $1.0000$
G1 $0.6469$ $1.0000$ $7.1\times 10^{-14}$ $3.5\times 10^{-13}$
G2 $0.0147$ $0.0734$ $1.1\times 10^{-15}$ $6.8\times 10^{-15}$
L1 $7\times 10^{-25} 2.2$$2\times 10^{-24}$ $3.3\times 10^{-150}$ $2.6\times 10^{-149}$
L2 $5\times 10^{-8} 1.1$$1\times 10^{-7}$ $5.2\times 10^{-76}$ $3.6\times 10^{-75}$
PI1 $0.5188$ $1.0000$ $0.8746$ $1.0000$
PI2 $0.0022$ $0.0131$ $4.6\times 10^{-11}$ $1.8\times 10^{-10}$
F-test for
Local Regression Model
pp-value $0.2979$ $5.3\times 10^{-14}$
Permutation Test for
Local Regression Model
pp-value $0.2379$ $0.0000$
* Use Holm-Bonferroni correction for multiple testing.
Table S6: Tests of the distribution difference of singularity scores in the mammary epithelial cell data. We verifiy the distribution difference using Spearman’s rank correlation tests between singularity scores and distances to cluster centers, F-tests and per- mutation tests for a local regression model (singularity scores regressed against locations). The results of all three tests confirm the distribution difference of singularity scores between small and large perplexities.
Embryo Singularity Score EMBEDR scDEED (elbow) scDEED DynamicViz
Perplexity 25 - 3 - 20
Neighborhood Preservation 0.5594 - 0.4955 - 0.5524
(a) Mouse embryonic stem cells (mESCs) differentiation data.
Brain Singularity Score EMBEDR scDEED (elbow) scDEED DynamicViz
Perplexity 95 145 10 145 10
Neighborhood Preservation 0.4108 0.4223 0.3575 0.4223 0.3575
(b) Mouse brain single-cell ATAC-seq data.
Mammary Singularity Score EMBEDR scDEED (elbow) scDEED DynamicViz
Perplexity 175 450 100 500 -
Neighborhood Preservation 0.6866 0.6937 0.6846 0.69387 -
(c) Mouse mammary epithelial cells data.
Table S7: Perplexities chosen by singularity score, EMBEDR, scDEED, and DynamicViz and the corresponding neighborhood preservation score. (a) Comparison of perplexity selection methods in mouse embryonic stem cell (mESC) differentiation data. EMBEDR fails to select a perplexity as we encountered an error when running the package, while scDEED’s second approach, which minimizes the number of dubious points, fails to yield a unique value. Our singular scores identify a perplexity that reduces sub-clusters while preserving the highest neighborhood continuity. (b) Comparison of perplexity selection methods in mouse brain single-cell ATAC-seq data. Our singular scores suggest a moderate perplexity that reduces sub-clusters, maintains high neighborhood preservation and retains fine-grained structure. (c) Comparison of perplexity selection methods in mouse mammary epithelial cells data. DynamicViz fails due to its lack of scalability for large datasets, as its bootstrap-based approach requires repeated execution of visualization algorithms. Among other methods, our singular scores suggest a moderate perplexity that ensures high neighborhood preservation while retaining fine-grained structure.
@@
Figure S2: Large perplexity lessens FI discontinuity. We use the same mixture data as in Figure 3 and run the standard t-SNE algorithm 6 times at perplexity 5, 20, 40, 60, 80, 100. a Contour plot shows the landscape of LOO loss $L({\mathbf{y}};{\mathbf{x}})$ for ${\mathbf{x}}=0.5({\mathbf{c}}_{1}+{\mathbf{c}}_{2})$ under different perplexities. The number of local minima in the loss decreases with higher perplexity, indicating that the FI discontinuity lessens under higher perplexity. b Trajectories of the embedding point of $\mathbf{x}$} are shown with different perplexities. Under small perplexities, numerous local minima cause an uneven trajectory of embedding points when we add $\mathbf{x}$} at evenly interpolated locations; while under larger perplexities, the trajectory is more smooth. This further suggests a reduction of FI discontinuity when the perplexity increases. c Embeddings with dichotomized singularity scores. Embedding points with high singularity scores decrease in number when increasing the perplexity.
@@
Figure S3: Comparing loss landscapes for single-cell data with different perplexities. We present LOO loss landscapes under small and large perplexities for mouse embryonic stem cells differentiation data (a-b) and mouse brain single-cell ATAC-seq data (c-d). We randomly choose an input ${\mathbf{x}}_{i}$ from the dataset and plot the landscape of the partial LOO loss $L_{i}({\mathbf{y}};{\mathbf{x}}_{i})$. We observe that more local minima emerge at random locations in the partial LOO loss landscape under a small perplexity than under a large perplexity, which supports our claim about the reduction of local minima and FI discontinuity under a large perplexity.
Figure S4: Comparing perturbation score with kernel PCA and one-class SVM for OOD detection. Using the dataset presented in Fig. 5, we compare the performance of the three methods. The perturbation score outperforms the others, achieving an average AUROC of 0.747 across the three selected clusters. In contrast, kernel PCA and one-class SVM obtain averaged AUROCs of 0.698 and 0.410, respectively.
Figure S5: Singularity scores inform selection of the perplexity parameter in the mouse brain single-cell ATAC-seq data. In the t-SNE visualizations of the mouse brain single-cell ATAC-seq data, multiple spurious sub-clusters are introduced at perplexity 5 (default by t-SNE algorithm) (a). High singularity scores appear in random locations, indicating the presence of such spurious structures and severe FI discontinuity. The perplexity chosen by the singularity score creates better visualization compared to default. Qualitatively, the cluster structure is more compact and the high singularity scores disappear except in the peripheries of the clusters. Quantitatively, the neighborhoods of most points are more faithfully embedded (higher neighbor-hood preservation) (b).
@@
Figure S6: Perturbation scores of CIFAR-10 image features with different choices of perturbation length. We vary the choice of the perturbation length $\lambda\in\{1,2,3\}$ in the calculation of perturbation scores (see Eqn. 6). We find that embedding points receiving high scores are consistent across different choices of $\lambda$. This suggests that perturbation scores are not sensitive to the choice of the perturbation length.
@@
Figure S7: Pre-screening points for calculating perturbations scores provides comparable results. We examine the validity of the pre-screening step by comparing the perturbation scores with the pre-screening step (visualized in left panels) and those without (visualized in right panels). We present the results of perturbation scores on three datasets—5-component Gaussian mixture data (a), Swiss roll data (b), and the deep learning feature data of CIFAR-10 (c). We find that the calculating perturbation scores with the pre-screening step still identifies most of the OI discontinuity locations, so pre-screening provides a faster and comparably reliable assessment of OI discontinuity.
@@
Figure S8: The approximation method 1&2 are accurate for calculating perturbation scores. We test the approximation quality on a dataset consisting of 5000 deep learning feature vectors obtained from CIFAR-10 images. a Scatter plot of exact perturbation scores v.s. perturbation scores by the approximation method 1. b Scatter plot of exact perturbation scores v.s. perturbation scores by approximation method 2. We find that the perturbation scores using both approximation methods are approximately equal to the exact perturbation scores for almost all the points.
@@
Figure S10: Comparing perplexity selection on the mouse embryonic cell differentiation data. a For mouse embryonic cell differentiation data, singularity score suggests selecting a perplexity of 25. DynamicViz chooses a perplexity of 20 by minimizing the variance score. By identifying the elbow point in the plot of dubious embedding points against perplexity, scDEED selects a perplexity of 3. However, it fails to select the optimal perplexity by minimizing the number of dubious points, as 26 out of 39 perplexity candidates reach the minimum number. EMBEDR is not applicable to the mouse embryonic cell differentiation data due to errors occurring for datasets smaller than 1,000. b The t-SNE embedding of perplexity 3 (chosen by scDEED) yields numerous spurious sub-clusters. The t-SNE embeddings of perplexity 20 (chosen by DynamicViz), 25 (chosen by singularity score), and 30 (default) are visually very similar.
Figure S11: Comparing perplexity selection on the mouse brain chromatin accessibility data. a For mouse brain chromatin accessibility data, singularity score suggests selecting a perplexity of 95. DynamicViz chooses a perplexity of 10. By identifying the elbow point in the plot of dubious embedding points against perplexity, scDEED also selects a perplexity of 10. And it selects a perplexity of 145 by minimizing the number of dubious points. EMBEDR also selects a perplexity of 145 by minimizing the median of the permutation test $p$-values, agreeing with its undesirable property of choosing the larger candidate perplexity. We used standard boxplots (center line, median; box limits, upper and lower quartiles; points, outliers). b The t-SNE embedding of perplexity 10 (chosen by DynamicViz and scDEED) yields numerous spurious sub-clusters. The t-SNE embeddings of perplexity 30 (default), 95 (chosen by singularity score), and 145 (chosen by scDEED and EMBEDR) are visually very similar. However, increasing perplexity excessively is not suggested, as it may cause clusters to merge and result in the loss of certain genuine local or microscopic structures.
Figure S12: Comparing perplexity selection on the mouse mammary epithelial cell dataset. a Singularity score suggests selecting a perplexity of 175. By identifying the elbow point in the plot of dubious embedding points against perplexity, scDEED selects a perplexity of 100. And it selects a perplexity of 500 by minimizing the number of dubious points. EMBEDR also selects a large perplexity of 450. We used standard boxplots (center line, median; box limits, upper and lower quartiles; points, outliers). DynamicViz lacks scalability for large datasets due to its bootstrap-based approach, which necessitates repeated execution of visualization algorithms. b The t-SNE embeddings of perplexity 100 (scDEED) and 175 (singularity score) are visually very similar. We can observe that clusters separated in perplexity 175 start to merge for larger perplexities of 450 and 500. Therefore, increasing perplexity excessively is not suggested.